# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 4.3 Parameter Interference and Sign Conflict Analysis
The parameter-level conflicts identified in the preceding discussion manifest through three distinct interference mechanisms that emerge when multiple task vectors are combined through element-wise operations. Systematic examination of these mechanisms is essential for developing principled mitigation strategies that preserve the compositional benefits of task arithmetic while minimizing performance degradation.
@@
Sign Conflicts represent perhaps the most pernicious form of parameter interference, occurring when task vectors from different fine-tuning objectives contain opposing directional updates for identical weight positions. Formally, for task vectors $\tau_{1},\tau_{2},\ldots,\tau_{n}$ derived from $n$ distinct fine-tuning procedures, a sign conflict at parameter position $j$ occurs when $\exists i,k:\text{sign}(\tau_{i}^{(j)})\neq\text{sign}(\tau_{k}^{(j)})$ and both $|\tau_{i}^{(j)}|,|\tau_{k}^{(j)}|>\epsilon$ for some threshold $\epsilon$. When such conflicts arise, naive averaging produces parameter values that represent a compromised intermediate state, potentially satisfying neither task’s requirements. Yadav et al. (2023) provided comprehensive empirical analysis demonstrating that sign conflicts correlate strongly with performance degradation in merged models, with certain layer types exhibiting systematically higher conflict rates than others. The analysis suggests that attention projection matrices and feed-forward intermediate layers may accumulate more severe conflicts, indicating that these components encode task-specific information that resists direct composition.
@@
Magnitude Disparities constitute a second critical interference mechanism, arising from the heterogeneous scaling of parameter updates across different fine-tuning procedures. Even when sign agreement exists, substantial differences in update magnitudes, quantified as $\max_{i}|\tau_{i}^{(j)}|/\min_{k}|\tau_{k}^{(j)}|$ for non-trivial updates, can cause certain tasks to dominate the merged representation disproportionately. This phenomenon is particularly pronounced when combining models fine-tuned with different learning rates, dataset sizes, or training durations (Yadav et al., 2023). The resulting merged model exhibits biased behavior favoring tasks whose fine-tuning produced larger magnitude updates, effectively drowning out the contributions of more conservatively trained variants. Subsequent investigations by Li et al. (2023a) demonstrate that real-world task vectors exhibit heavy-tailed magnitude distributions with significant inter-task variance.
@@
Redundant Parameter Modifications introduce a subtler but equally important source of interference, wherein low-magnitude updates that do not encode meaningful task-specific knowledge accumulate noise during the merging process. These redundant parameters arise from optimization dynamics that produce non-zero gradients even for weights irrelevant to the target task, creating spurious modifications that dilute the signal-to-noise ratio of the merged representation. Yu et al. (2023) formalized this observation by analyzing the effective rank and information content of task vectors, finding that a substantial fraction of parameter updates contribute minimally to task performance while actively degrading merge quality through interference with truly informative modifications.
@@
The interaction between these interference mechanisms produces complex, often non-additive effects that resist simple analytical characterization. (Li et al., 2023a) introduced statistical frameworks for analyzing the joint distribution of task vector parameters, revealing that interference severity depends critically on the correlation structure between task vectors rather than their marginal statistics alone. Visualization techniques including parameter-wise conflict heatmaps and layer-wise interference scores have emerged as diagnostic tools for identifying problematic regions before merging. This comprehensive characterization of interference phenomena, spanning sign conflicts, magnitude disparities, and redundant modifications, directly motivates the sparsification-enhanced approaches that selectively retain informative parameters while eliminating conflict-inducing modifications.
@@ 4.4 Sparsification-Enhanced Merging Methods
Sparsification-enhanced methods address these interference phenomena by exploiting redundancy in task vectors: strategic parameter elimination reduces conflicts while preserving task-critical information. These approaches transform dense task vectors into sparse representations with improved composability.
The seminal Trim-Elect-Sign (TIES) mechanism introduced a three-stage pipeline that systematically addresses each category of interference. Given task vectors $\{\tau_{i}\}_{i=1}^{n}$ derived from $n$ fine-tuned models, the trimming stage first eliminates parameters with magnitudes below a threshold $\lambda$, producing sparse vectors $\tilde{\tau}_{i}$ where $\tilde{\tau}_{i,j}=\tau_{i,j}\cdot\mathbf{1}$. The election stage then resolves sign conflicts through majority voting, computing a consensus sign mask $\gamma_{j}=\text{sgn}(\sum_{i}\tilde{\tau}_{i,j})$ that determines the dominant direction for each parameter across all task vectors. Finally, the sign alignment stage retains only those parameter updates consistent with the elected sign, yielding conflict-free task vectors suitable for aggregation.
@@
0: Task vectors $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$, trim ratio $k$, pretrained weights $\theta_{\text{pre}}$
@@
Yadav et al. (2023) demonstrated that this mechanism achieves substantial performance improvements over naive task arithmetic, particularly when merging more than three task vectors where sign conflicts proliferate.
@@
Complementing the deterministic TIES approach, the DARE (Drop And REscale) framework (Yu et al., 2023) introduced probabilistic sparsification as an alternative conflict resolution strategy. Rather than employing magnitude-based pruning, DARE applies random dropout to task vector parameters with probability $p$, followed by rescaling of the retained parameters by factor 1/(1p)1/(1-p) to preserve the expected contribution. Mathematically, the sparsified task vector becomes $\hat{\tau}_{i}=\frac{1}{1-p}\cdot\tau_{i}\odot m_{i}$ where $m_{i}$ is a binary mask with $m_{i,j}\sim\text{Bernoulli}(1-p)$. This approach draws theoretical justification from the observation that fine-tuning induces primarily low-rank modifications to pretrained weights, similar to findings in network pruning research (Frankle and Carbin, 2019), implying substantial redundancy that can be exploited through random sampling without catastrophic information loss. Recent extensions such as DELLA-Merging (Deep et al., 2024) further refine this approach through magnitude-based sampling (Deep et al., 2024) strategies. The stochastic nature of DARE provides an implicit ensemble effect when multiple merging trials are conducted, though practical deployments typically employ a single sparsification pass.
@@
Building upon these foundational methods, importance-aware sparsification has been proposed that leverages second-order information to guide parameter retention decisions. Rather than treating all parameters equivalently during the trimming process, such methods compute importance scores based on the Fisher information or gradient magnitude accumulated during fine-tuning, preferentially retaining parameters whose modification contributed most significantly to task adaptation. This approach addresses a key limitation of magnitude-based pruning, wherein parameters with small absolute values may nonetheless encode critical task-specific information.
@@
Subsequent work extended the sparsification paradigm by introducing conflict-awareness into the retention decision process. Unlike TIES, which considers sign conflicts only after independent trimming of each task vector, conflict-aware methods jointly optimize sparsification masks across all source models to minimize expected interference while maximizing information retention. These methods employ a balanced sparsification criterion that normalizes contribution magnitudes across tasks, preventing dominant task vectors from overwhelming weaker contributors during the merge.
@@
Further refinements have incorporated distributional statistics of task vector parameters into sparsification decisions. Recognizing that optimal trimming thresholds vary across layers and parameter groups, these methods adaptively calibrate sparsification intensity based on local magnitude distributions, achieving more precise conflict resolution than global threshold approaches. While these sparsification techniques have proven effective at mitigating interference, they nonetheless operate within the constraints of direct weight-space manipulation, motivating alternative formulations that reconsider the mathematical foundations of task vector composition itself.
@@ 4.5 Advanced Task Vector Manipulation and Recent Extensions
Motivated by these limitations of direct weight-space manipulation, recent methodological advances have extended the task arithmetic framework toward more principled mathematical reformulations that address fundamental limitations of standard parameter arithmetic. A particularly influential direction involves reconceptualizing task vector operations within the tangent space of neural network parameters, motivated by the observation that standard weight-space arithmetic implicitly assumes local linearity that may not hold for large parameter displacements.
@@
Research in this area introduced a linearized formulation that operates on task vectors projected into the tangent space centered at the pretrained model weights (Tang et al., 2024c). Given a pretrained model with parameters $\theta_{0}$ and a fine-tuned model $\theta_{t}$, the standard task vector $\tau_{t}=\theta_{t}-\theta_{0}$ captures parameter displacements without accounting for the nonlinear relationship between weights and model behavior. The tangent space approach instead considers the linearized model $f_{\text{lin}}(x;\theta)=f(x;\theta_{0})+\nabla_{\theta}f(x;\theta_{0})^{\top}(\theta-\theta_{0})$, which provides a first-order approximation of how parameter changes affect model outputs. Within this framework, task vector arithmetic operates on the linearized representations, yielding improved compositional properties when combining multiple task-specific adaptations. Empirically, tangent space task arithmetic demonstrates superior performance on model editing benchmarks, particularly for negation operations where the standard formulation often produces unstable results.
@@
The theoretical motivation underlying tangent space methods connects to the broader question of why task arithmetic succeeds despite the highly nonlinear nature of neural network loss landscapes. The linearization effectively constrains operations to a regime where the superposition principle approximately holds, addressing scenarios where direct weight interpolation violates the implicit linearity assumptions of standard task arithmetic. This reformulation proves especially valuable for combining task vectors with large magnitudes or when source models have undergone extensive fine-tuning that displaces parameters significantly from the pretrained initialization.
@@
Complementary to these mathematical reformulations, adaptive rank-based approaches have emerged as powerful techniques for optimizing the representational structure of task vectors. Recent work proposed methods that adaptively determine optimal rank configurations for low-rank approximations of task vectors across different layers and modules. Unlike uniform rank reduction strategies, these approaches recognize that different network components encode task-specific information with varying degrees of redundancy, allocating representational capacity accordingly. The method formulates rank selection as an optimization problem that balances compression efficiency against information preservation, leveraging singular value decomposition to identify the effective dimensionality of task-specific modifications.
The integration of adaptive rank methods with existing sparsification techniques yields synergistic benefits. While sparsification operates by zeroing individual parameters, rank-based approaches restructure the entire task vector representation, capturing different aspects of redundancy in fine-tuned models. Table 4 summarizes these methods. This complementarity enables hybrid strategies that first reduce task vector rank to eliminate correlated modifications, then apply sparsification to address residual conflicts among the compressed representations. Li et al. (2023a) provided comprehensive empirical analysis demonstrating that such combined approaches consistently outperform either technique in isolation across diverse task combinations.
@@
From a practical implementation perspective, these advanced manipulation methods have been progressively incorporated into community toolkits that facilitate reproducible model merging experimentation. The modular architecture of modern merging frameworks enables researchers to compose tangent space projections, adaptive rank reduction, and conflict-aware sparsification into customized merging pipelines tailored to specific application requirements. This ecosystem integration significantly lowers barriers to adoption, allowing practitioners to leverage sophisticated mathematical formulations without implementing complex algorithms from scratch. Despite their increasing sophistication, however, all task vector manipulation techniques discussed thus far share a common assumption: that merging should ultimately produce a unified parameter set. An alternative paradigm instead preserves distinct expert components while enabling flexible, input-dependent utilization of specialized capabilities.
Table 4: Comparison of task vector arithmetic and sparsification-enhanced methods for model merging.
@@
DARE Drop and Rescale Randomly drops delta parameters and rescales remaining ones to preserve expectations Stochastic nature may introduce variance 2024
CABS Conflict-Aware Balanced Sparsification Explicitly identifies and eliminates conflicting parameters while maintaining balance Limited to identical architectures 2024
@@
Task vector arithmetic provides a powerful conceptual framework for model editing, enabling intuitive operations such as knowledge addition and negation, while sparsification-enhanced methods like TIES-Merging and DARE address critical challenges of parameter interference and sign conflicts by selectively pruning redundant or conflicting parameters before aggregation. These approaches operate primarily at the individual parameter level, treating model weights as independent quantities to be manipulated through algebraic operations. While this section focused on parameter-space manipulations and sparsity-driven conflict resolution, the next section examines structured and information-guided merging approaches that move beyond element-wise operations to exploit architectural organization through Mixture-of-Experts (Shazeer et al., 2017) compositions, leverage activation statistics for principled merging decisions, and employ evolutionary search to automatically discover optimal merging configurations across the full design space.

### 30 Mar 2026 v2

@@ 4.3 Parameter Interference and Sign Conflict Analysis
Parameter-level conflicts identified in the preceding discussion manifest through three distinct interference mechanisms that emerge when multiple task vectors are combined through element-wise operations. Detailed examination of these mechanisms is essential for developing principled mitigation strategies that preserve the compositional benefits of task arithmetic while minimizing performance degradation.
@@
Sign Conflicts represent perhaps the most pernicious form of parameter interference, occurring when task vectors from different fine-tuning objectives contain opposing directional updates for identical weight positions. Formally, for task vectors $\tau_{1},\tau_{2},\ldots,\tau_{n}$ derived from $n$ distinct fine-tuning procedures, a sign conflict at parameter position $j$ occurs when $\exists i,k:\text{sign}(\tau_{i}^{(j)})\neq\text{sign}(\tau_{k}^{(j)})$ and both $|\tau_{i}^{(j)}|,|\tau_{k}^{(j)}|>\epsilon$ for some threshold $\epsilon$. When such conflicts arise, naive averaging produces parameter values that represent a compromised intermediate state, potentially satisfying neither task’s requirements. Yadav et al. (2023) provided detailed empirical analysis showing that sign conflicts correlate strongly with performance degradation in merged models, motivating their three-step trim-elect-merge procedure for conflict resolution.
@@
Magnitude Disparities constitute a second critical interference mechanism, arising from the heterogeneous scaling of parameter updates across different fine-tuning procedures. Even when sign agreement exists, large differences in update magnitudes, quantified as $\max_{i}|\tau_{i}^{(j)}|/\min_{k}|\tau_{k}^{(j)}|$ for non-trivial updates, can cause certain tasks to dominate the merged representation disproportionately. This phenomenon is particularly pronounced when combining models fine-tuned with different learning rates, dataset sizes, or training durations (Yadav et al., 2023). The resulting merged model exhibits biased behavior favoring tasks whose fine-tuning produced larger magnitude updates, effectively drowning out the contributions of more conservatively trained variants. Empirical investigations reveal that real-world task vectors exhibit heavy-tailed magnitude distributions with substantial inter-task variance.
@@
Redundant Parameter Modifications introduce a subtler but equally important source of interference, wherein low-magnitude updates that do not encode meaningful task-specific knowledge accumulate noise during the merging process. These redundant parameters arise from optimization dynamics that produce non-zero gradients even for weights irrelevant to the target task, creating spurious modifications that dilute the signal-to-noise ratio of the merged representation. Yu et al. (2024) formalized this observation by showing that delta parameters from fine-tuning are extremely sparse, with the vast majority contributing minimally to task performance while actively degrading merge quality through interference with truly informative modifications.
@@
Interactions between these interference mechanisms produce complex, often non-additive effects that resist simple analytical characterization. Statistical frameworks for analyzing the joint distribution of task vector parameters reveal that interference severity depends strongly on the correlation structure between task vectors rather than their marginal statistics alone. Visualization techniques including parameter-wise conflict heatmaps and layer-wise interference scores have become diagnostic tools for identifying problematic regions before merging. This careful characterization of interference phenomena, spanning sign conflicts, magnitude disparities, and redundant modifications, directly motivates the sparsification-enhanced approaches that selectively retain informative parameters while eliminating conflict-inducing modifications.
@@ 4.4 Sparsification-Enhanced Merging Methods
Sparsification-enhanced methods mitigate these interference phenomena by exploiting redundancy in task vectors, as strategic parameter elimination reduces conflicts while preserving task-critical information. These approaches transform dense task vectors into sparse representations with improved composability.
The seminal Trim-Elect-Sign (TIES) mechanism introduced a three-stage pipeline that systematically resolves each category of interference. Given task vectors $\{\tau_{i}\}_{i=1}^{n}$ derived from $n$ fine-tuned models, the trimming stage retains only the top-kk% of parameters by magnitude, producing sparse vectors $\tilde{\tau}_{i}$ where all other entries are set to zero. The election stage then resolves sign conflicts through majority voting, computing a consensus sign mask $\gamma_{j}=\text{sign}(\sum_{i}\tilde{\tau}_{i,j})$ that determines the dominant direction for each parameter across all task vectors. Finally, the sign alignment stage retains only those parameter updates consistent with the elected sign, yielding conflict-free task vectors suitable for aggregation.
@@
0: Task vectors $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$, trim ratio $k$, scaling coefficient $\lambda$, pretrained weights $\theta_{\text{pre}}$
@@
Yadav et al. (2023) showed that this mechanism achieves substantial performance improvements over naive task arithmetic, particularly when merging more than three task vectors where sign conflicts proliferate.
@@
Complementing the deterministic TIES approach, the DARE (Drop And REscale) framework (Yu et al., 2024) introduced probabilistic sparsification as an alternative conflict resolution strategy. Rather than using magnitude-based pruning, DARE applies random dropout to task vector parameters with probability $p$, followed by rescaling of the retained parameters by factor 1/(1p)1/(1-p) to preserve the expected contribution. Mathematically, the sparsified task vector becomes $\hat{\tau}_{i}=\frac{1}{1-p}\cdot\tau_{i}\odot m_{i}$ where $m_{i}$ is a binary mask with $m_{i,j}\sim\text{Bernoulli}(1-p)$. This approach draws theoretical justification from the observation that fine-tuning induces primarily low-rank modifications to pretrained weights, similar to findings in network pruning research (Frankle and Carbin, 2019), implying considerable redundancy that can be exploited through random sampling without catastrophic information loss. Recent extensions such as DELLA-Merging (Deep et al., 2024) further refine this approach through magnitude-based sampling strategies. The stochastic nature of DARE provides an implicit ensemble effect when multiple merging trials are conducted, though practical deployments typically use a single sparsification pass.
@@
Building upon these foundational methods, importance-aware sparsification has been proposed that uses second-order information to guide parameter retention decisions. Rather than treating all parameters equivalently during the trimming process, such methods compute importance scores based on the Fisher information or gradient magnitude accumulated during fine-tuning, preferentially retaining parameters whose modification contributed most directly to task adaptation. This approach overcomes a key limitation of magnitude-based pruning, wherein parameters with small absolute values may nonetheless encode critical task-specific information.
@@
Subsequent work extended the sparsification approach by introducing conflict-awareness into the retention decision process. Unlike TIES, which considers sign conflicts only after independent trimming of each task vector, conflict-aware methods jointly optimize sparsification masks across all source models to minimize expected interference while maximizing information retention. CABS (Conflict-Aware Balanced Sparsification) (Yang et al., 2025e) introduces explicit conflict detection and balanced sparsification that normalizes contribution magnitudes across tasks, preventing dominant task vectors from overwhelming weaker contributors during the merge. Concrete Subspace Learning (Tang et al., 2023) takes an alternative approach by learning continuous relaxations of discrete sparsification masks through concrete distributions, allowing gradient-based optimization of interference elimination. Sens-Merging (Liu et al., 2025a) proposes sensitivity-guided parameter balancing that uses gradient sensitivity analysis to determine which parameters most require conflict resolution.
@@
Further refinements have incorporated distributional statistics of task vector parameters into sparsification decisions. DPPA (Zhu et al., 2024b) introduces a dual-stage method combining dynamic magnitude-based pruning with partition-based amplification (rescaling) to better preserve important parameters while resolving conflicts during merging. ImPart (Yang et al., 2025b) proposes importance-aware delta-sparsification that jointly considers compression and merging objectives, striking improved trade-offs between model size and merged performance. Recognizing that optimal trimming thresholds vary across layers and parameter groups, these methods adaptively calibrate sparsification intensity based on local magnitude distributions, producing more precise conflict resolution than global threshold approaches. While these sparsification techniques have proven effective at mitigating interference, they nonetheless operate within the constraints of direct weight-space manipulation, motivating alternative formulations that reconsider the mathematical foundations of task vector composition itself.
@@ 4.5 Advanced Task Vector Manipulation and Recent Extensions
Motivated by these limitations of direct weight-space manipulation, recent methodological advances have extended the task arithmetic framework toward more principled mathematical reformulations that tackle core limitations of standard parameter arithmetic. A particularly influential direction involves reconceptualizing task vector operations within the tangent space of neural network parameters, motivated by the observation that standard weight-space arithmetic implicitly assumes local linearity that may not hold for large parameter displacements.
@@
Research in this area introduced a partial linearization formulation for parameter-efficient adapters (Tang et al., 2024d), which linearizes only the adapter modules and applies task arithmetic over the linearized representations. Given a pretrained model with parameters $\theta_{\text{pre}}$ and a fine-tuned model $\theta_{t}$, the standard task vector $\tau_{t}=\theta_{t}-\theta_{\text{pre}}$ captures parameter displacements without accounting for the nonlinear relationship between weights and model behavior. The tangent space approach instead considers the linearized model $f_{\text{lin}}(x;\theta)=f(x;\theta_{\text{pre}})+\nabla_{\theta}f(x;\theta_{\text{pre}})^{\top}(\theta-\theta_{\text{pre}})$, which provides a first-order approximation of how parameter changes affect model outputs. Within this framework, task vector arithmetic operates on the linearized representations, yielding improved compositional properties when combining multiple task-specific adaptations. Empirically, tangent space task arithmetic yields superior performance on model editing benchmarks, particularly for negation operations where the standard formulation often produces unstable results. Recent theoretical analysis by Li et al. (2025b) provide formal guarantees for when task vectors are provably effective for model editing in nonlinear Transformers, establishing generalization bounds that help practitioners predict merging success.
@@
The theoretical appeal of tangent space methods stems from their explicit accommodation of neural network nonlinearity. The linearization constrains operations to a regime where the superposition principle approximately holds, handling scenarios where direct weight interpolation violates the implicit linearity assumptions of standard task arithmetic. This reformulation proves especially valuable for combining task vectors with large magnitudes or when source models have undergone extensive fine-tuning that moves parameters appreciably from the pretrained initialization.
@@
Complementary to these mathematical reformulations, adaptive rank-based approaches optimize the representational structure of task vectors. Methods in this direction adaptively determine optimal rank configurations for low-rank approximations of task vectors across different layers and modules. STAR (Spectral Truncation and Rescale) (Lee et al., 2025c) uses spectral analysis to identify and preserve the most informative components of task vectors while truncating noise. LoRE-Merging (Liu et al., 2025b) explores low-rank estimation specifically designed for large language model merging, delivering effective compression of task vectors with minimal capability loss. No Task Left Behind (Marczak et al., 2025) proposes isotropic model merging that operates in common and task-specific subspaces to ensure that all constituent tasks are adequately represented. Superpose Task-specific Features (Qiu et al., 2025a) introduces a feature superposition framework that supports more effective combination of task-specific information. Unlike uniform rank reduction strategies, these approaches recognize that different network components encode task-specific information with varying degrees of redundancy, allocating representational capacity accordingly. The method formulates rank selection as an optimization problem that balances compression efficiency against information preservation, using singular value decomposition to identify the effective dimensionality of task-specific modifications.
Integrating adaptive rank methods with existing sparsification techniques yields synergistic benefits. While sparsification operates by zeroing individual parameters, rank-based approaches restructure the entire task vector representation, capturing different aspects of redundancy in fine-tuned models. Layer-Aware Task Arithmetic (Chen et al., 2025f) disentangles task-specific knowledge from instruction-following capabilities at the layer level, allowing more targeted composition. Adaptive Task Vectors (Kang et al., 2025a) dynamically generate input-conditioned task vectors using a small language model, enabling query-specific adaptation without fixed demonstration sets. Decom-Renorm-Merge (Chaichana et al., 2025) proposes performing merging in a decomposed and renormalized space, showing that operating in the right mathematical space markedly improves multi-tasking performance. Optimal Brain Iterative Merging (Wang et al., 2025d) draws inspiration from second-order pruning methods to iteratively mitigate interference during the merging process. Table 4 summarizes these methods. This complementarity enables hybrid strategies that first reduce task vector rank to eliminate correlated modifications, then apply sparsification to resolve residual conflicts among the compressed representations.
@@
From a practical standpoint, community toolkits such as MergeKit (Goddard et al., 2024) have progressively incorporated these advanced manipulation methods, offering modular pipelines that compose tangent space projections, adaptive rank reduction, and conflict-aware sparsification into reproducible merging workflows. Despite their increasing complexity, however, all task vector manipulation techniques discussed thus far share a common assumption, namely that merging should produce a unified parameter set. An alternative approach instead preserves distinct expert components while allowing flexible, input-dependent use of specialized capabilities.
Table 4: Task vector arithmetic and sparsification-enhanced methods for model merging. Each method addresses a specific aspect of parameter interference: sign conflicts (TIES-Merging), redundant updates (DARE), conflict-aware balancing (CABS), or tangent space reformulation. The “Interference Handling” column indicates whether the method explicitly resolves parameter conflicts between task vectors.
@@
DARE Drop and Rescale Randomly drops delta parameters and rescales remaining ones to preserve expectations Stochastic nature may introduce variance 2023
CABS Conflict-Aware Balanced Sparsification Explicitly identifies and eliminates conflicting parameters while maintaining balance Limited to identical architectures 2025
@@
Task vector arithmetic provides a powerful conceptual basis for model editing, supporting intuitive operations such as knowledge addition and negation, while sparsification-enhanced methods like TIES-Merging and DARE tackle critical challenges of parameter interference and sign conflicts by selectively pruning redundant or conflicting parameters before aggregation. Beyond these core methods, recent extensions reveal several converging trends that illuminate the field’s trajectory.
A first trend is the refinement of sparsity and pruning criteria. While TIES-Merging and DARE establish that sparsification improves merging, follow-up work shows that the choice of which parameters to prune matters as much as the pruning rate. LEWIS (Chopra et al., 2025) proposes layer-wise sparsity for training-free guided merging, recognizing that different layers require different pruning intensities, while AdaRank (Lee et al., 2025a) replaces element-wise sparsity with adaptive rank pruning that operates on the singular value structure of task vectors. This progression from uniform to adaptive sparsification mirrors the broader shift observed in weight-space averaging methods.
A second, closely related trend moves conflict resolution from parameter space to representation space. Representation Surgery (Yang et al., 2024c) mitigates multi-task interference directly at the representation level, Subspace-Boosted Merging (Skorobogat et al., 2025) exploits subspace decomposition, and Purifying Task Vectors (An et al., 2025) operates in knowledge-aware subspaces. The key insight is that parameter-level sign conflicts may be symptoms of deeper representational incompatibilities; by resolving interference in the representation domain, these methods can address root causes rather than surface-level symptoms.
Third, several methods introduce explicit stability constraints to prevent catastrophic forgetting during merging. Task Arithmetic in Trust Region (Sun et al., 2025b) constrains merged parameters to remain within trust regions around the pretrained initialization, LINES (Wang et al., 2025b) prevents forgetting via post-training layer scaling, and Adaptive Weight Disentanglement (Xiong et al., 2024) separates task-specific from shared components. These approaches trade merging flexibility for stronger preservation guarantees, a tradeoff that becomes increasingly important as the number of merged tasks grows.
Fourth, differentiable and optimization-based formulations recast merging as a continuous optimization problem rather than a discrete recipe. DAM (Gauthier-Caron et al., 2024) provides a differentiable path from simple averaging to full model composition, while adaptive projective gradient descent (Wei et al., 2025c) enables gradient-based coefficient learning. By making the merging process end-to-end differentiable, these methods can jointly optimize merging coefficients, but at the cost of requiring validation data, a departure from the data-free approach that characterizes most task arithmetic methods.
Fifth, cross-model transfer methods relax the strict shared-initialization requirement. Update Your Transformer (Rinaldi et al., 2025a) supports re-basin of task vectors across pretrained model versions, Gradient-Sign Masking (Rinaldi et al., 2025b) enables task vector transport, and Decomposing Task Vectors (Damirchi et al., 2025) supports refined model editing across model families. These methods represent an important step toward broader applicability, though they typically incur additional alignment costs compared to same-initialization merging.
Finally, novel applications show the versatility of the task vector abstraction beyond multi-task combination. Personality Vector (Sun et al., 2025a) extends task arithmetic to modulate personality traits of LLMs through continuous-strength control and multi-trait composition. Expert Merging (Zhang et al., 2025b) proposes unsupervised expert alignment via hidden-state and logit matching with importance-guided layer chunking. Efficient Compositional Multi-tasking (Bohdal et al., 2025) introduces learnable calibration for merging task adapters when a single input must be processed for multiple tasks simultaneously on edge devices. Additional contributions include SE-Merging (Chen et al., 2025g), MetaGPT (Zhou et al., 2024b), MoD (Dang and Ngo, 2024), distribution-aware sparse fusion (Lin et al., 2026), and ParamDelta (Cao et al., 2025). While this section focused on parameter-space manipulations and sparsity-driven conflict resolution, the next section examines structured and information-guided merging approaches that move beyond element-wise operations to exploit architectural organization through Mixture-of-Experts (Shazeer et al., 2017) compositions, use activation statistics for guided merging decisions, and apply evolutionary search to automatically discover optimal merging configurations across the full design space.